# Quantum Generative Models for Image Generation: Insights from MNIST and MedMNIST

### v1

@@ Abstract
Research on quantum generative models is currently in its early exploratory stages, with very few established methodologies. In this paper, we propose a novel hybrid quantum generative model based on variational quantum circuits for image generation tasks, introducing innovative noise techniques specifically tailored for quantum computation. Our approach utilizes two distinctive noise strategies: quantum-generated noise inherent to quantum circuits, and a newly developed noise scheduling method, applying different noise levels strategically across time steps during the training process. Experiments conducted on MNIST and MedMNIST datasets demonstrate that our hybrid quantum model, combined with these specialized noise techniques, achieves promising results, suggesting improved generative performance compared to baseline quantum generative approaches. This exploratory work lays a critical foundation and opens new avenues for advancing quantum generative modeling research.
@@ IV-B Diffusion Process
IV-C Improved U-Net Architecture and Timestep Embedding
@@ IV-E Training Strategy and EMA
V Experiments and Results
V-A Datasets and Experimental Setup
@@ V-D Results on MedMNIST
VI Discussion And Conclusion
In this work, we proposed a variational quantum circuit-based generative model for image synthesis, implemented using the PennyLane hybrid quantum-classical framework. Through empirical evaluation on both MNIST and PathMNIST datasets, we explored the strengths and limitations of quantum diffusion-based generative approaches compared to classical baselines.
Our experiments reveal that the quantum model performs competitively on simpler grayscale datasets like MNIST, achieving higher average SSIM and lower FID in some cases, particularly in capturing fine-grained digit structures. These results suggest that the quantum architecture can efficiently encode structural patterns within compact circuits, even under current hardware constraints. However, the performance advantage does not fully carry over to more complex and high-resolution color datasets. On PathMNIST, the quantum model achieved a better FID score, indicating improved alignment with the global data distribution, but suffered from significantly lower SSIM values, reflecting a lack of fine-grained spatial fidelity in generated images.
This discrepancy highlights a key limitation: while quantum models are capable of capturing global distributional properties, they may struggle with preserving detailed local features, particularly in high-dimensional, color-rich image domains. This suggests that current quantum circuit designs may be more suited to learning abstract or low-dimensional representations rather than detailed spatial structures, which are crucial in domains such as medical imaging.

### v2

@@ Abstract
Quantum generative models offer a promising new direction in machine learning by leveraging quantum circuits to enhance data generation capabilities. In this study, we propose a hybrid quantum-classical image generation framework that integrates variational quantum circuits into a diffusion-based model. To improve training dynamics and generation quality, we introduce two novel noise strategies: intrinsic quantum-generated noise and a tailored noise scheduling mechanism. Our method is built upon a lightweight U-Net architecture, with the quantum layer embedded in the bottleneck module to isolate its effect. We evaluate our model on MNIST and MedMNIST datasets to examine its feasibility and performance. Notably, our results reveal that under limited data conditions (fewer than 100 training images), the quantum-enhanced model generates images with higher perceptual quality and distributional similarity than its classical counterpart using the same architecture. While the quantum model shows advantages on grayscale data such as MNIST, its performance is more nuanced on complex, color-rich datasets like PathMNIST. These findings highlight both the potential and current limitations of quantum generative models and lay the groundwork for future developments in low-resource and biomedical image generation.
@@ III-B Diffusion Process
III-C Improved U-Net Architecture and Timestep Embedding
@@ III-E Training Strategy and EMA
IV Experiments and Results
IV-A Datasets and Experimental Setup
@@ IV-D Results on MedMNIST
V Discussion And Conclusion
In this work, we proposed a variational quantum circuit-based generative diffusion model for image synthesis, implemented using the PennyLane hybrid quantum-classical framework. Through empirical evaluation on both MNIST and PathMNIST datasets, we explored the strengths and limitations of quantum diffusion-based generative approaches compared to classical baselines.
Our experiments reveal that the quantum model performs competitively on simpler grayscale datasets like MNIST, achieving higher average SSIM and lower FID in some cases, particularly in capturing fine-grained digit structures. Notably, under low-data regimes—when trained with fewer than 100 samples—the quantum model consistently outperformed the classical counterpart using the same architecture, indicating superior generalization ability and robustness in data-scarce settings. These findings suggest that quantum circuits can effectively encode structural patterns within compact representations, even under current hardware constraints.
However, this performance advantage does not fully carry over to more complex and high-resolution color datasets. On PathMNIST, the quantum model achieved a better FID score, indicating improved alignment with the global data distribution, but suffered from significantly lower SSIM values, reflecting a lack of fine-grained spatial fidelity in generated images.
This discrepancy highlights a key limitation: while quantum models are capable of capturing global distributional properties, they may struggle with preserving detailed local features—especially in high-dimensional, color-rich image domains. This suggests that current quantum circuit designs may be better suited for learning abstract or low-dimensional representations rather than detailed spatial structures, which are crucial in domains such as medical imaging.